# Wenbin Wang, Wenjie Xu, Colin N. Jones

### v1

@@ IV Performance Analysis
We use the expected distance to $u^{*}$, i.e., $\mathbb{E}[\|u_{k}-u^{*}\|^{2}]$, as the error metric, as it is commonly employed in the literature for online zeroth-order optimization [22, 27]. Our results indicate that for a fixed time step $k^{\prime}$, $\mathbb{E}[\|u_{k}-u^{*}\|^{2}]$ is bounded by $\mathbb{E}[\|u_{k^{\prime}}-u^{*}\|^{2}]$, scaled by an exponentially decaying factor, along with a constant term $\mathcal{O}(\mu,\mu^{k^{\prime}},\delta)$, which is the function of $k^{\prime}$. At equilibrium, $\mathbb{E}[\|u_{k^{\prime}}-u^{*}\|^{2}]$ vanishes, leaving the left-hand side bounded solely by $\mathcal{O}(\mu,\mu^{k^{\prime}},\delta)$. Since $\mu^{k^{\prime}}$ decreases as $k^{\prime}$ increases for $\mu<1$, the steady-state error is eventually fully characterized by $\mathcal{O}(\mu,\delta)$.
@@ V Numerical simulation
To show that the controller (8) is capable of identifying $u^{*}$, we conduct several simulations with an LTI system,
@@
where $A\in\mathbb{R}^{n_{x}\times n_{x}}$, $B\in\mathbb{R}^{n_{x}\times n_{u}}$ are the system matrices. This system has a steady-state input-output map $H=(I-A)^{-1}B$ and is pre-stabilized by a lower-level controller, ensuring that the spectral radius of $A$ is less than one. Such systems are commonly found in various applications, including building control [28], and power systems [24].
@@ V-A A simple example
To analyze the impact of plant transient on algorithm performance, we first consider a quadratic problem
@@
where the solution can be written as $u=H^{-1}x_{\text{ref}}$. We assign $A=\begin{bmatrix}c&1\\ 0&c\end{bmatrix}$, $B$ to be the identity matrix and $x_{\text{ref}}=[100,100]^{\top}$. We switch $c$ from $c=0.1$ to $c=0.7$ to represent plants with different decay rates. We set $\eta=0.1$ and $\delta=0.5$ in the simulation. Note that a quadric function does not have a bounded gradient. Hence, Assumption 2 is not satisfied. The result can be found in Fig. 1.
@@
In Fig. 1, the relative error $\|x_{k}-x_{\text{ref}}\|/\|x_{\text{ref}}\|$ in logarithmic scale is plotted for $c=0.1$ and $c=0.7$. The solid line represents the mean value over 20 simulations, while the transparent area is the standard deviation. The orange line represents the algebraic plant, i.e., assuming $H$ is known and directly sampling $\bm{1}_{u_{k}+\delta v_{k}\succ u_{k-1}+\delta v_{k-1}}$, and the blue line represents (8). When at the steady state, both (8) and its steady-state counterpart return the result with the same level of precision. However, for a slower system ($c\!=\!0.7$), (8) has a larger overshoot and a higher variance in steady state.
@@ V-B Thermal comfort optimization
Next, we consider a thermal comfort optimization problem where a building is modeled via an LTI system with 13 states [28]. We model an occupant’s utility on thermal comfort with the well-known Predictive Mean Vote (PMV) model [3]. The PMV model is usually quantified via Predicted Percentage of Dissatisfied (PPD), which is a nonconvex function with respect to room temperature. We generate the preference feedback based on this model, aiming to explore different room temperatures to minimize PPD using controller (8). For other inputs required by the PMV model, we assume the occupant is typing and wearing sweatpants, T-shirt and Shoes or sandals (default setting in the PMV model).
In Fig. 2, the mean indoor temperature (blue solid line) is plotted over 20 runs of simulations. The black horizontal line represents the ground truth. With careful tuning of $\eta$ and $\delta$, we found that (8) can track the optimal point effectively without large overshoot, which demonstrates its potential in learning a human’s utility in real-world applications.
@@ VI Conclusion
In this work, we designed a human-aware controller that utilizes preference feedback to optimize utility in a dynamic system. We derived an explicit upper bound on the error introduced by approximating the steady-state input-state map using real-state measurements and random perturbations. We analyzed the impact of both the system decay rate and smoothing parameters on the stability and optimality of the closed-loop system. Numerical experiments on thermal comfort optimization demonstrate its potential for solving real-world problems. Further directions include extending the theoretical framework to alternative user models, e.g., noise-free, and exploring real-world applications in areas like product design and chemical selection.

### v2

@@ IV Performance Analysis
We adopt the expected distance to $u^{*}$, i.e., $\mathbb{E}[\|u_{k}-u^{*}\|^{2}]$, as the error metric, which is commonly used in the literature on online zeroth-order optimization [22, 26]. Our results indicate that, for any fixed time step $k^{\prime}$, $\mathbb{E}[\|u_{k}-u^{*}\|^{2}]$ is bounded by $\mathbb{E}[\|u_{k^{\prime}}-u^{*}\|^{2}]$, scaled by an exponentially decaying factor, and a constant term of order $\mathcal{O}(\mu,\mu^{k^{\prime}},\delta)$ that depends on $k^{\prime}$. In steady state, $\mathbb{E}[\|u_{k^{\prime}}-u^{*}\|^{2}]$ vanishes, so that the bound reduces to $\mathcal{O}(\mu,\mu^{k^{\prime}},\delta)$. Since $\mu^{k^{\prime}}$ decreases as $k^{\prime}$ increases for $\mu<1$, the steady-state error is eventually fully characterized by $\mathcal{O}(\mu,\delta)$.
@@ V Numerical simulation
To demonstrate that the controller (8) is capable of identifying $u^{*}$, we perform numerical simulations on an LTI system defined by
@@
where $A\in\mathbb{R}^{n_{x}\times n_{x}}$, $B\in\mathbb{R}^{n_{x}\times n_{u}}$ are the system matrices. This system has an invertible steady-state input-output map $H=(I-A)^{-1}B$ and is pre-stabilized by a lower-level controller, ensuring that the spectral radius of $A$ is less than one. Such systems frequently arise in applications, including building control [27], and power systems [24].
@@ V-A A simple example
To analyze the impact of plant transients on the algorithm’s performance, we first consider the quadratic problem
@@
where the solution is $u=H^{-1}x_{\text{ref}}$. We set $A\!=\!\begin{bmatrix}c&1\\ 0&c\end{bmatrix}$, $B$ to be the identity matrix and $x_{\text{ref}}=[100,100]^{\top}$. The parameter $c$ varies between $c=0.1$ and $c=0.7$ to represent plants with different decay rates. The simulation parameters are chosen as $\eta=0.1$ and $\delta=0.5$. The results are shown in Fig. 1.
@@
In Fig. 1, the relative error $\|x_{k}-x_{\text{ref}}\|/\|x_{\text{ref}}\|$ on a logarithmic scale is shown for $c=0.1$ and $c=0.7$. The solid line represents the mean value over 20 simulations, while the shaded region indicates one standard deviation. The orange line corresponds to the algebraic plant, which assumes $H$ is known and samples $\bm{1}_{u_{k}+\delta v_{k}\succ u_{k-1}+\delta v_{k-1}}$ directly, whereas the blue line represents (8). At the steady state, both (8) and its steady-state counterpart achieve a comparable level of accuracy. However, for the slower system (c=0.7c\!=\!0.7), (8) exhibits a larger overshoot and higher steady-state variance.
@@ V-B Thermal comfort optimization
Next, we consider a thermal comfort optimization problem, in which a building is represented by an LTI system with 13 states [27]. An occupant’s thermal comfort utility is represented with the well-known Predictive Mean Vote (PMV) model [3]. The PMV output is typically expressed in terms of the Predicted Percentage of Dissatisfied (PPD) index, which is a nonconvex function of room temperature. Preference feedback is generated based on this model, with the goal of identifying the room temperature that minimizes PPD using the controller (8). For other PMV parameters, we assume the occupant is typing, wearing sweatpants, T-shirt and shoes or sandals (default setting in the PMV model).
In Fig. 2, the mean indoor temperature with (8) over 20 simulations (blue line) is plotted, while the orange line represents the case with a noise-free user model, i.e., $\bm{1}_{(x_{k+1},u_{k}+\delta v_{k})\succ(x_{k},u_{k-1}+\delta v_{k-1})}=\text{sign}(\Phi(x_{k},u_{k-1}+\delta v_{k-1})-\Phi(x_{k+1},u_{k}+\delta v_{k}))$. The shaded region represents one standard deviation, and the black horizontal line represents the true optimal temperature. With careful tuning of $\eta$ and $\delta$, controller (8) can track the optimal point effectively without large overshoot. which demonstrates its potential for learning a human’s utility in real-world applications. Meanwhile, the algorithm exhibits a higher convergence rate when there is no noise in the preference feedback. Quantifying its closed-loop behavior from a theoretical perspective remains an interesting direction for future work.
@@ VI Conclusion
In this work, we developed a human-aware controller that utilizes preference feedback to optimize utility while accounting for system transients. We derived an explicit upper bound on the error introduced by approximating the steady-state input-state map using the real-time state measurements. We analyzed the impact of both the system decay rate and smoothing parameters on the stability and optimality of the closed-loop system. Numerical experiments on a thermal comfort optimization task demonstrate its potential for solving real-world problems. Further research directions include extending the theoretical framework to alternative user models (e.g., noise-free model) and exploring real-world applications such as product design and chemical selection.